# A representational framework for learning and encoding structurally enriched trajectories in complex agent environments

### v2

@@ 1 Introduction
Autonomous systems perceive their environment, take actions, and adapt their behaviour through experience. One key paradigm in this field is Reinforcement learning (RL) (Sutton et al.,, 1998), which enables agents to learn from interaction with their environment by using reward signals to guide decision-making. Unlike supervised learning, which relies on labelled data, or unsupervised learning, which identifies patterns in data, RL allows agents to develop policies that maximise cumulative rewards over time. This approach has led to significant advancements in Artificial Intelligence (AI), for instance in game-playing AI (Silver et al.,, 2017), robotics (Levine et al.,, 2018), and autonomous systems (Kiran et al.,, 2021). Despite these successes, RL agents face fundamental challenges when transferring knowledge from simpler environments to complex, dynamic ones requiring strong generalisation and knowledge transfer.
Representation learning plays a crucial role in addressing these challenges by developing structured, transferable encodings of an agent’s environment (Bengio et al.,, 2013; Lesort et al.,, 2018). Various approaches have been proposed, for example, hierarchical RL (Pateria et al.,, 2021) introduces structured abstractions for multi-step decision-making but remains constrained by predefined reward structures (Sutton et al.,, 1999; Bacon et al.,, 2017). Goal-conditioned RL enhances planning by leveraging future states as learning targets, yet still depends on explicitly defined objectives (Kaelbling,, 1993; Schaul et al.,, 2015). Disentanglement techniques refine state representations by separating controllable and uncontrollable factors, improving sample efficiency (Higgins et al.,, 2018).
Although these proposals have been instrumental in enhancing policy learning, improving sample efficiency, and facilitating task-specific generalisation, moving beyond compression methods towards enriching structural dependencies between states and actions is a critical step for more robust and general-purpose intelligence (Nie et al.,, 2017).
@@
Encoding isolated interactions between objects, however, is not enough; agents need to also learn how these interactions influence future decisions and outcomes. To model this effectively, it is essential to incorporate affordances, which describe the potential interactions available to an agent based on both its capabilities and the environment (Gibson,, 2014). An affordance is a property of an object in relation to others and from the perspective of an agent that prompts a specific use or interaction. For example, a chair may afford sitting to a human but not to a large animal, such as an elephant. Structured representations of affordances enable an agent to predict the effects of its interactions, providing a foundation for adaptive motor planning and effective task generalisation (Schack and Ritter,, 2013), also enable agents to anticipate how changes in context impact success or failure, reducing the risk of misgeneralisation (Khazatsky et al.,, 2021).
Despite all the advantages that representing affordances would introduce, they alone would only apply to a given state. That is, affordances do not track how interactions unfold sequentially. Tracking how interactions unfold sequentially requires maintaining a structured history of transitions, a trajectory. By retaining this structured trajectory rather than relying solely on the Markov assumption that states encapsulate all relevant history, disregarding transitions (Niv,, 2019), agents will no longer be compelled to make decisions based only on their current state and will be able to learn the dependencies that shape long-term task execution, enabling more adaptive and informed decision-making. Encoding such sequences of interactions and their distant effects has been shown to be essential for AI systems to adapt to novel scenarios (Eysenbach et al.,, 2022), enabling the agent to capture long-term dependencies and avoid inefficient learning or goal misgeneralisation (Di Langosco et al.,, 2022).
@@
These developments present an opportunity to expand the representation of trajectories with hierarchical structures. Unlike traditional representations that treat states and actions as independent transitions, graph structures naturally encode dependencies across multiple levels of abstraction, preserving how interactions evolve over time. As the need for modelling complex dependencies increases, Graph Neural Networks (GNNs) have become a key tool for learning from structured data, enabling AI systems to capture rich relational information. Heterogeneous Graph Neural Networks (HGNNs) (Wang et al.,, 2021) extend this capability by incorporating multiple node and edge types, allowing a more detailed representation of diverse interactions in dynamic environments. Still, despite their merits, most existing graph-based approaches focus on encoding full graphs rather than leveraging subgraphs with shared components. This limits their ability to model localised structures while capturing commonalities across different subgraphs. By treating entire graphs as independent entities, these methods often fail to exploit the reusability of nodes, making generalisation more challenging. A subgraph-based approach, where fundamental components are shared, offers a more effective way to represent task dynamics and facilitate knowledge transfer.
@@ 2 Related work
A core challenge in reinforcement learning (RL) is generalisation. While agents can master specific tasks, they often struggle to transfer knowledge to new situations. This limitation stems largely from how they represent their experiences. Early approaches focused on learning compressed state representations (Wang et al.,, 2021), aiming to distil high-dimensional observations into compact vectors. However, while efficient, these static snapshots often fail to generalize well, as not all aspects of a state are transferable to new tasks .
A more powerful approach is to learn from trajectories, the full sequence of states and actions leading to a result. Yet, the first and most basic information gleaned from a trajectory is its outcome (Williams,, 1992). Knowing whether a sequence of actions led to success or failure is a crucial learning signal. But relying solely on outcomes is insufficient for transferring knowledge. For instance, the outcome of travelling to a destination can be achieved by either riding a bike or driving a car. While the outcome is the same, the skills required are entirely different (Wilson and Collins,, 2019). Knowing only that both methods were ’successful’ provides no transferable knowledge about how to perform either task. This emphasises that understanding the trajectory, the way in which a task is completed, is crucial for adapting and transferring skills.
Furthermore, even learning from trajectories may not suffice to capture the nuances of information composition and organisation. Classically, they are defined as ”flat” sequences of state-action pairs, a format that lacks deep structural information. This overlooks the rich internal structure that humans intuitively grasp. A key strength of human problem-solving is the ability to develop hierarchical and causal models of our interactions, allowing us to recognize structural similarities and adapt our strategies dynamically (Schack and Ritter,, 2013; Morris et al.,, 2022). The key challenge for AI, therefore, is not just to record trajectories, but to learn to represent their internal structure. This section reviews the evolution of attempts to solve this problem, from early temporal abstractions to modern, heterogeneous graph-based methods.
@@ 2.1 Early Trajectory Representations: From Flat Sequences to Temporal Abstractions
Methods like the Outcome-Driven Actor-Critic (ODAC) trajectory-based prediction models, and memory-based RL frameworks such as Sequential Episodic Control (SEC) all adopt simple, shallow sequences of state-action pairs (Rudner et al.,, 2021; Haines et al.,, 2018; Freire et al.,, 2025), focusing on linear dependencies rather than hierarchical structures (Sutton et al.,, 1998).
@@ 2.2 The Shift to Structured Representations: Skills and Latent Spaces
To better capture a trajectory’s internal structure, subsequent research focused on learning representations of entire trajectories. The Self-Consistent Trajectory Autoencoder (SeCTAR) model framed the problem as one of representation learning, building continuous latent spaces of trajectories to represent structured behaviours (Co-Reyes et al.,, 2018). While a significant step, SeCTAR’s latent space does not explicitly segment trajectories into transferable skills, limiting its use for compositional learning.
In contrast, algorithms like Constructing Skill Trees (CST) automatically segment demonstration trajectories into reusable skills (Konidaris et al.,, 2011). CST enabled the decomposition of long tasks into meaningful sub-policies. However, CST’s reliance on human-provided demonstrations means it cannot autonomously discover skills through its own exploration.
@@ 2.3 Relational Learning: The Emergence of Graph-Based Approaches
Graph-based approaches provide a natural way to represent this structured knowledge. Heterogeneous Graph Neural Networks (HGNNs) are particularly well-suited, as they can encode multi-level interactions between different types of nodes (e.g., objects, agents, affordances) using mechanisms like meta-paths (Wang et al.,, 2021). Some research has explored causal graphs to model structured transitions rather than mere correlations, though these often rely on semi-supervised learning, making them unsuitable for pure RL settings (Jin et al.,, 2024).
@@ 2.4 The State of the Art: Heterogeneous and Self-Supervised Graph Models
Recent advances in HGNNs have focused on self-supervised learning. Heterogeneous Co-contrastive Learning (HeCo) uses two complementary views (network schema and meta-path) to learn rich node embeddings for static graphs (Wang et al.,, 2021). Graph Transformer Co-contrastive Learning (GTC) extends this with a Transformer-based architecture to better capture both local and global dependencies (Sun et al.,, 2025).
@@ 3.1.2 Affordances
Following (Catarau-Cotutiu et al.,, 2023), an affordance is conceptualised as the learnt relationships between an (effect, reward) pair and a (concept-object, interaction) tuple. When an agent applies an action, it produces an interaction, which in turn generates an effect (a state transition) and an associated reward, both of which are stored as an affordance. This structured representation is designed to improve generalise across tasks by predicting how different interactions influence trajectories across varying contexts.
Formally, an affordance $A$ encapsulates an agent’s action, the current state, and the resulting trajectory (Catarau-Cotutiu et al.,, 2023):
@@ 3.1.5 Structured Data Sources: CREATE and MiniGrid
Our primary environment is Construction, Reuse, and Extension of Actions through Transfer and Exploration (CREATE) (Jain et al.,, 2020). CREATE is a diverse and challenging platform designed to study action generalisation, providing tasks that involve dynamic interactions with various objects and tools in a continuous physics-based simulation. It allows us to explore generalisation by varying task setups, forcing agents to adapt strategies and reuse knowledge from past experiences.
@@ 3.2.1 SET Definition Through Random Exploration
In RL, an episode consists of a complete sequence of interactions from an initial state to termination, encapsulating the trajectory an agent follows within a given task. To extract and formalise SETs (for a detailed definition, see Section 3.1.3) we employ an exploration strategy within the CREATE environment (Jain et al.,, 2020). In this setup, agents engage with tasks such as pushing a ball to a goal or placing it inside a bucket, generating diverse trajectories. To maximise variability in the trajectory space and better capture adaptive patterns and task-relevant structures, RL agents operate under a random policy rather than optimised for specific rewards.
@@ 3.2.2 Building the Hierarchical Memory Graph
To effectively model SETs, we structured episodic data into hierarchical heterogeneous graphs, where the feature extraction functions $\Phi:V\to\mathbb{R}^{d}$ (see Section 3.1.3) assigned high-dimensional embeddings to nodes representing objects, interactions, and states. The nodes and their different relations are organised into hierarchical heterogeneous graphs with three distinct levels of abstraction: (1) a base layer for objects and interactions, where objects are extracted using the SAM model (Kirillov et al.,, 2023) and interaction dynamics are encoded via a ConvLSTM; (2) a mid-layer for states and affordances, capturing the transition dynamics and relationships between interactions and their effects; and (3) a trajectory layer for task-level abstractions, representing the cumulative impact of interactions and sequences (see Fig. 2). The selection of these specific models was based on a comparative analysis to ensure robustness and generalisability, as explained in our component analysis in Section 4.2. This hierarchical organisation is stored in a graph database and is not simply a stratification of layers but a structured representation where lower levels influence higher-level abstractions. Objects and actions influence affordances, which in turn shape state transitions and contribute to the formation of SETs. This dependency across levels enables agents to reason over task structure, capturing both local dynamics and high-level task generalisation in complex environments. Formally, the hierarchical memory graph G representing a SET consists of:
@@
Figure 3: Representation of two objects interacting. Interactions are characterized not only by their execution , but also by the relationships they share with the objects involved. On the right, the different parameters of an Interaction are shown such as the type, and the vector representation. Representation obtained from the Neo4j (Lal,, 2015) graph database interface.
@@
    Objects within each episode are extracted using the SAM (Segment Anything Model) object detection model Kirillov et al., (2023), which identifies and classifies entities in the environment. A process of checking whether the object exists already is carried out using a cosine similarity function.
@@
    Figure 4: Visualisation of the interaction encoder’s ability to learn semantically meaningful representations. This scatter plot shows K-means clusters of interaction embeddings, categorized by the physical properties of the object involved (’Bouncy’ vs. ’Fixed’). The clear separation between the two groups demonstrates that the encoder has successfully learned to distinguish between these two functionally critical classes of interactions from visual data alone, a key prerequisite for building the SET hierarchy.
@@ 3.3.1 Graph Encoder Architecture Based on Heterogeneous Co-contrastive Learning (HeCo)
Graph encoders rely on message passing, a core operation in Graph Neural Networks (GNNs) (Scarselli et al.,, 2008), where each node iteratively aggregates information from its neighbours to refine its representation (Gilmer et al.,, 2017). In its general form, message passing is defined as:
@@ 4.4 Comparison on CREATE environment
In Jain et al., (2020), action representations were evaluated by encoding agent actions across episodes and analysing their separability within the embedding space. Their results highlighted effective clustering of actions, particularly for distinguishing tasks requiring distinct agent behaviours.
In contrast, SETLE shifted the focus to trajectory representations, encoding not only the actions but also the intermediate states, transitions, and affordances that cumulatively define the task’s success or failure. The comparison reveals that action representations, as emphasised by Jain et al., (2020), excel in precision for short-term predictions, such as determining the next best action in a sequence. SETLE, however, takes a broader approach by also learning interaction representations using a ConvLSTM architecture and then integrating them with state, affordance, and object information to enable comprehensive SET representation learning. By combining detailed interaction-level encoding with higher-level abstractions of task trajectories, SETLE provides a unified framework that supports both immediate decision-making and generalisation across tasks. This integration highlights the complementary nature of interaction and trajectory representations, with SETLE excelling in task-level analysis while maintaining robust interaction representation capabilities.

### v3

@@ 1 Introduction
Autonomous systems perceive their environment, take actions, and adapt their behaviour through experience. One key paradigm in this field is Reinforcement learning (RL) (sutton1998reinforcement), which enables agents to learn from interaction with their environment by using reward signals to guide decision-making. Unlike supervised learning, which relies on labelled data, or unsupervised learning, which identifies patterns in data, RL allows agents to develop policies that maximise cumulative rewards over time. This approach has led to significant advancements in Artificial Intelligence (AI), for instance in game-playing AI (silver2017mastering), robotics (levine2018learning), and autonomous systems (kiran2021deep). Despite these successes, RL agents face fundamental challenges when transferring knowledge from simpler environments to complex, dynamic ones requiring strong generalisation and knowledge transfer.
Representation learning plays a crucial role in addressing these challenges by developing structured, transferable encodings of an agent’s environment (bengio2013representation; lesort2018state). Various approaches have been proposed, for example, hierarchical RL (pateria2021hierarchical) introduces structured abstractions for multi-step decision-making but remains constrained by predefined reward structures (sutton1999between; bacon2017option). Goal-conditioned RL enhances planning by leveraging future states as learning targets, yet still depends on explicitly defined objectives (kaelbling1993learning; schaul2015universal). Disentanglement techniques refine state representations by separating controllable and uncontrollable factors, improving sample efficiency (higgins2018towards).
Although these proposals have been instrumental in enhancing policy learning, improving sample efficiency, and facilitating task-specific generalisation, moving beyond compression methods towards enriching structural dependencies between states and actions is a critical step for more robust and general-purpose intelligence (nie2017).
@@
Encoding isolated interactions between objects, however, is not enough; agents need to also learn how these interactions influence future decisions and outcomes. To model this effectively, it is essential to incorporate affordances, which describe the potential interactions available to an agent based on both its capabilities and the environment (gibson1977theory). An affordance is a property of an object in relation to others and from the perspective of an agent that prompts a specific use or interaction. For example, a chair may afford sitting to a human but not to a large animal, such as an elephant. Structured representations of affordances enable an agent to predict the effects of its interactions, providing a foundation for adaptive motor planning and effective task generalisation (schack2013representation), also enable agents to anticipate how changes in context impact success or failure, reducing the risk of misgeneralisation (khazatsky2021affordances).
Despite all the advantages that representing affordances would introduce, they alone would only apply to a given state. That is, affordances do not track how interactions unfold sequentially. Tracking how interactions unfold sequentially requires maintaining a structured history of transitions, a trajectory. By retaining this structured trajectory rather than relying solely on the Markov assumption that states encapsulate all relevant history, disregarding transitions (niv2019learning), agents will no longer be compelled to make decisions based only on their current state and will be able to learn the dependencies that shape long-term task execution, enabling more adaptive and informed decision-making. Encoding such sequences of interactions and their distant effects has been shown to be essential for AI systems to adapt to novel scenarios (eysenbach2022contrastive), enabling the agent to capture long-term dependencies and avoid inefficient learning or goal misgeneralisation (di2022goal).
@@
These developments present an opportunity to expand the representation of trajectories with hierarchical structures. Unlike traditional representations that treat states and actions as independent transitions, graph structures naturally encode dependencies across multiple levels of abstraction, preserving how interactions evolve over time. As the need for modelling complex dependencies increases, Graph Neural Networks (GNNs) have become a key tool for learning from structured data, enabling AI systems to capture rich relational information. Heterogeneous Graph Neural Networks (HGNNs) (wang2021selfsupervised) extend this capability by incorporating multiple node and edge types, allowing a more detailed representation of diverse interactions in dynamic environments. Still, despite their merits, most existing graph-based approaches focus on encoding full graphs rather than leveraging subgraphs with shared components. This limits their ability to model localised structures while capturing commonalities across different subgraphs. By treating entire graphs as independent entities, these methods often fail to exploit the reusability of nodes, making generalisation more challenging. A subgraph-based approach, where fundamental components are shared, offers a more effective way to represent task dynamics and facilitate knowledge transfer.
@@ 2 Related work
A core challenge in reinforcement learning (RL) is generalisation. While agents can master specific tasks, they often struggle to transfer knowledge to new situations. This limitation stems largely from how they represent their experiences. Early approaches focused on learning compressed state representations (wang2021selfsupervised), aiming to distil high-dimensional observations into compact vectors. However, while efficient, these static snapshots often fail to generalize well, as not all aspects of a state are transferable to new tasks .
A more powerful approach is to learn from trajectories, the full sequence of states and actions leading to a result. Yet, the first and most basic information gleaned from a trajectory is its outcome (williams1992reinforce). Knowing whether a sequence of actions led to success or failure is a crucial learning signal. But relying solely on outcomes is insufficient for transferring knowledge. For instance, the outcome of travelling to a destination can be achieved by either riding a bike or driving a car. While the outcome is the same, the skills required are entirely different (collins2021). Knowing only that both methods were ’successful’ provides no transferable knowledge about how to perform either task. This emphasises that understanding the trajectory, the way in which a task is completed, is crucial for adapting and transferring skills.
Furthermore, even learning from trajectories may not suffice to capture the nuances of information composition and organisation. Classically, they are defined as ”flat” sequences of state-action pairs, a format that lacks deep structural information. This overlooks the rich internal structure that humans intuitively grasp. A key strength of human problem-solving is the ability to develop hierarchical and causal models of our interactions, allowing us to recognize structural similarities and adapt our strategies dynamically (schack2013representation; morris2022neural). The key challenge for AI, therefore, is not just to record trajectories, but to learn to represent their internal structure. This section reviews the evolution of attempts to solve this problem, from early temporal abstractions to modern, heterogeneous graph-based methods.
@@ 2.1 Early Trajectory Representations: From Flat Sequences to Temporal Abstractions
Methods like the Outcome-Driven Actor-Critic (ODAC) trajectory-based prediction models, and memory-based RL frameworks such as Sequential Episodic Control (SEC) all adopt simple, shallow sequences of state-action pairs (rudner2021outcome; haines2018outcome; freire2024sequential), focusing on linear dependencies rather than hierarchical structures (sutton1998reinforcement).
@@ 2.2 The Shift to Structured Representations: Skills and Latent Spaces
To better capture a trajectory’s internal structure, subsequent research focused on learning representations of entire trajectories. The Self-Consistent Trajectory Autoencoder (SeCTAR) model framed the problem as one of representation learning, building continuous latent spaces of trajectories to represent structured behaviours (co-reyes2018sectar). While a significant step, SeCTAR’s latent space does not explicitly segment trajectories into transferable skills, limiting its use for compositional learning.
In contrast, algorithms like Constructing Skill Trees (CST) automatically segment demonstration trajectories into reusable skills (konidaris2011cst). CST enabled the decomposition of long tasks into meaningful sub-policies. However, CST’s reliance on human-provided demonstrations means it cannot autonomously discover skills through its own exploration.
@@ 2.3 Relational Learning: The Emergence of Graph-Based Approaches
Graph-based approaches provide a natural way to represent this structured knowledge. Heterogeneous Graph Neural Networks (HGNNs) are particularly well-suited, as they can encode multi-level interactions between different types of nodes (e.g., objects, agents, affordances) using mechanisms like meta-paths (wang2021selfsupervised). Some research has explored causal graphs to model structured transitions rather than mere correlations, though these often rely on semi-supervised learning, making them unsuitable for pure RL settings (jin2024causal).
@@ 2.4 The State of the Art: Heterogeneous and Self-Supervised Graph Models
Recent advances in HGNNs have focused on self-supervised learning. Heterogeneous Co-contrastive Learning (HeCo) uses two complementary views (network schema and meta-path) to learn rich node embeddings for static graphs (wang2021selfsupervised). Graph Transformer Co-contrastive Learning (GTC) extends this with a Transformer-based architecture to better capture both local and global dependencies (sun2025gtc).
@@ 3.1.2 Affordances
Following (catarau2023aigenc), an affordance is conceptualised as the learnt relationships between an (effect, reward) pair and a (concept-object, interaction) tuple. When an agent applies an action, it produces an interaction, which in turn generates an effect (a state transition) and an associated reward, both of which are stored as an affordance. This structured representation is designed to improve generalise across tasks by predicting how different interactions influence trajectories across varying contexts.
Formally, an affordance $A$ encapsulates an agent’s action, the current state, and the resulting trajectory (catarau2023aigenc):
@@ 3.1.5 Structured Data Sources: CREATE and MiniGrid
Our primary environment is Construction, Reuse, and Extension of Actions through Transfer and Exploration (CREATE) (jain2020generalization). CREATE is a diverse and challenging platform designed to study action generalisation, providing tasks that involve dynamic interactions with various objects and tools in a continuous physics-based simulation. It allows us to explore generalisation by varying task setups, forcing agents to adapt strategies and reuse knowledge from past experiences.
@@ 3.2.1 SET Definition Through Random Exploration
In RL, an episode consists of a complete sequence of interactions from an initial state to termination, encapsulating the trajectory an agent follows within a given task. To extract and formalise SETs (for a detailed definition, see Section 3.1.3) we employ an exploration strategy within the CREATE environment (jain2020generalization). In this setup, agents engage with tasks such as pushing a ball to a goal or placing it inside a bucket, generating diverse trajectories. To maximise variability in the trajectory space and better capture adaptive patterns and task-relevant structures, RL agents operate under a random policy rather than optimised for specific rewards.
@@ 3.2.2 Building the Hierarchical Memory Graph
To effectively model SETs, we structured episodic data into hierarchical heterogeneous graphs, where the feature extraction functions $\Phi:V\to\mathbb{R}^{d}$ (see Section 3.1.3) assigned high-dimensional embeddings to nodes representing objects, interactions, and states. The nodes and their different relations are organised into hierarchical heterogeneous graphs with three distinct levels of abstraction: (1) a base layer for objects and interactions, where objects are extracted using the SAM model (kirillov2023segment) and interaction dynamics are encoded via a ConvLSTM; (2) a mid-layer for states and affordances, capturing the transition dynamics and relationships between interactions and their effects; and (3) a trajectory layer for task-level abstractions, representing the cumulative impact of interactions and sequences (see Fig. 2). The selection of these specific models was based on a comparative analysis to ensure robustness and generalisability, as explained in our component analysis in Section 4.2. This hierarchical organisation is stored in a graph database and is not simply a stratification of layers but a structured representation where lower levels influence higher-level abstractions. Objects and actions influence affordances, which in turn shape state transitions and contribute to the formation of SETs. This dependency across levels enables agents to reason over task structure, capturing both local dynamics and high-level task generalisation in complex environments. Formally, the hierarchical memory graph G representing a SET consists of:
@@
Figure 3: Representation of two objects interacting. Interactions are characterized not only by their execution , but also by the relationships they share with the objects involved. On the right, the different parameters of an Interaction are shown such as the type, and the vector representation. Representation obtained from the Neo4j (lal2015neo4j) graph database interface.
@@
    Objects within each episode are extracted using the SAM (Segment Anything Model) object detection model kirillov2023segment, which identifies and classifies entities in the environment. A process of checking whether the object exists already is carried out using a cosine similarity function.
@@
    Figure 4: Visualisation of the interaction encoder’s ability to learn semantically meaningful representations. PCA projection of interaction embeddings with overlaid K-means cluster centroids. Light points represent individual samples, while larger markers denote centroids, with color indicating physical property (Bouncy vs. Fixed) and marker shape representing object type. A clear separation emerges along PC1 (38.3% variance), showing that physical interaction type is the dominant latent factor learned by the encoder. The encoder has successfully learned to distinguish between these two functionally critical classes of interactions from visual data alone, a key prerequisite for building the SET hierarchy.
@@ 3.3.1 Graph Encoder Architecture Based on Heterogeneous Co-contrastive Learning (HeCo)
Graph encoders rely on message passing, a core operation in Graph Neural Networks (GNNs) (scarselli2009gnn), where each node iteratively aggregates information from its neighbours to refine its representation (gilmer2017neural). In its general form, message passing is defined as:
@@ 4.4 Comparison on CREATE environment
In jain2020generalization, action representations were evaluated by encoding agent actions across episodes and analysing their separability within the embedding space. Their results highlighted effective clustering of actions, particularly for distinguishing tasks requiring distinct agent behaviours.
In contrast, SETLE shifted the focus to trajectory representations, encoding not only the actions but also the intermediate states, transitions, and affordances that cumulatively define the task’s success or failure. The comparison reveals that action representations, as emphasised by jain2020generalization, excel in precision for short-term predictions, such as determining the next best action in a sequence. SETLE, however, takes a broader approach by also learning interaction representations using a ConvLSTM architecture and then integrating them with state, affordance, and object information to enable comprehensive SET representation learning. By combining detailed interaction-level encoding with higher-level abstractions of task trajectories, SETLE provides a unified framework that supports both immediate decision-making and generalisation across tasks. This integration highlights the complementary nature of interaction and trajectory representations, with SETLE excelling in task-level analysis while maintaining robust interaction representation capabilities.